HanAkoSolo (jamierox4ev3r):

can anyone solve: (a^2-a-2)*(a^2+2a+1) ???

OpenStudy (anonymous):

(A+1)(A^3-3a-2)

OpenStudy (anonymous):

It cant be factored any further utalizing real numbers.

HanAkoSolo (jamierox4ev3r):

so it can be factored with imaginary numbers?

OpenStudy (anonymous):

I believe so. You can attempt but imaginary numbers arnt my strong suit sorry.

HanAkoSolo (jamierox4ev3r):

mine either thats cool

OpenStudy (anonymous):

(a+1) (a-2) * (a+1) (a+1) = 0 (a+1)^3 (a-2) =0 (a+1)^3 = 0 and (a-2)= 0 so a = -1 and a = 2

HanAkoSolo (jamierox4ev3r):

hey thanks camila! i would give u guys both medals.. but mike was first